# Tracing the bottom electroweak dipole operators at future lepton colliders

### v1

@@ Abstract
While often omitted in the SMEFT analyses of electroweak measurements, the electroweak dipole operators of the bottom quark have been found to be important in some cases, and are also related to processes involving the top quark. In this paper, we further investigate their effects, focusing on the measurements of the $e^{+}e^{-}\to b\bar{b}$ process at a future lepton collider. Their linear contributions are suppressed by the bottom mass due to the helicity flip in the interference term with the SM amplitude, leading to a nontrivial interplay between the linear and the quadratic contributions. With two runs at the Z-pole and 240 GeV, the effects of CP-even dipole coefficients can be well separated from the modifications of the SM $Zb\bar{b}$ couplings, while an additional run (e.g. at 360 GeV) is useful for lifting a nontrivial second best-fit point due to the quadratic contributions. The dipole operators also offer an alternative explanation to the long-standing $A^{0,b}_{\rm FB}$ discrepancy at LEP, which can be tested at a future lepton collider.
@@ I Introduction
The Standard Model (SM) of particle physics has been tested by collider experiments for several decades. In the electroweak (EW) sector, the SM predictions are in good agreements with the precision measurements carried out at the Large Electron-Positron collider (LEP) and Stanford Linear Collider (SLC) collider experiments ALEPH:2005ab ; ALEPH:2013dgf , as well as the ones at the Large Hadron Collider (LHC). Higgs measurements have been carried out in multiple channels at the LHC, and a good precision has already been achieved in some channels with the accumulated data ATLAS:2022vkf ; CMS:2022dwd . Several future lepton collider projects have also been proposed, including CEPC CEPCStudyGroup:2018ghi ; CEPCPhysicsStudyGroup:2022uwl , FCC-ee FCC:2018evy ; Bernardi:2022hny , ILC ILCInternationalDevelopmentTeam:2022izu , C3 Bai:2021rdg , CLIC CLICdp:2018cto ; CLIC:2018fvx , as well as a possible high-energy muon collider Aime:2022flm ; Accettura:2023ked . These future projects hold the potential to significantly enhance the current measurement precision of Higgs and EW observables, providing a even better probe of the SM and the possible new physics beyond it.
Assuming the scale of the potential new physics is significantly larger than the EW scale, the Standard Model Effective Field Theory (SMEFT) provides a systematic and model independent parameterization of the new physics effects, where deviations from the SM predictions (from new physics effects) are given by the Wilson coefficients of higher dimensional operators,
@@
where $d$ is the dimension of the operator and $\Lambda$ parameterizes the scale of the potential underlying new physics. The dimension-six operators are expected to constitute the leading effects to the experimental observables (besides the dimension-5 operator which generates Majorana neutrino masses), and have been thoroughly studied. Many global analyses have been performed, with both the current measurements and the projected future ones Falkowski:2015jaa ; Durieux:2017rsg ; Barklow:2017suo ; Durieux:2018tev ; Durieux:2018ggn ; Ellis:2018gqa ; DeBlas:2019qco ; Durieux:2019rbz ; Ellis:2020unq ; Bruggisser:2022rhb ; Liu:2022vgo ; deBlas:2022ofj ; Ethier:2021bye ; Brivio:2022hrb ; Bartocci:2023nvp ; Allwicher:2023shc ; Grunwald:2023nli ; Garosi:2023yxg ; Elmer:2023wtr ; Kassabov:2023hbm ; Bellafronte:2023amz ; Celada:2024mcf ; Asteriadis:2024qim . In many cases, a good overall precision can be achieved despite the large number of free parameters — a clear demonstration of the usefulness of the global SMEFT framework.
@@
While the top dipole operators are considered in many SMEFT analyses at high-energy colliders, the dipole operators of other fermions are usually omitted, and for good reasons. [^1] On the other hand, see Refs. daSilvaAlmeida:2019cbr ; Wen:2024cfu on studies of light-quark dipoles at colliders and Ref. Wen:2023xxc on probing electron dipole operators with transverse spin asymmetry. First, the SM amplitude and the amplitude with one dipole insertion have different helicity structures, and a mass insertion is needed to obtain an interference term. For light fermions, this means that the new physics effects are strongly suppressed by fermion mass at the $1/\Lambda^{2}$ level and mainly given by the dipole-squared terms, which are at the $1/\Lambda^{4}$ level. Second, low energy electric or magnetic dipole measurements provide very stringent constraints on some of the dipole operators (especially the electron and muon ones Muong-2:2006rrc ; ACME:2018yjb ; Muong-2:2008ebm ; Muong-2:2021ojo ) which are far beyond the reaches of collider experiments. However, for the second heaviest fermion, the bottom quark, the above two reasons are not fully applicable, and the effects of dipole operators are indeed found to be important in some cases. For instance, in Ref. Liu:2022vgo , it was found that the linear contributions of bottom dipole in EW observables, while suppressed by the bottom mass, tend to be more important than the loop effects of top operators, and the former should not be ignored if the latter is included in the analysis. On the other hand, Ref. Durieux:2019rbz studied the quadratic contributions of bottom dipole operators (while setting the bottom mass to zero) and found their effects to be important in a global framework. Together, these two studies suggest that both the linear and the quadratic contributions of the bottom dipole operators could be important, and an analysis including the full contributions of bottom dipole operators is needed to understand their impacts.
@@
In this paper, we conduct a further investigation into the effects of bottom dipole operators in the $e^{+}e^{-}\to b\bar{b}$ process, with their full contributions (both linear and quadratic) taken into account. We focus on the scenario of a future circular lepton collider, with a tera-Z program that significantly improves the EW measurements at LEP/SLD, as well as runs at both 240240240\,240GeV and around the $t\bar{t}$ threshold (around 360360360\,360GeV). As a first step towards a more comprehensive global analysis, in this study we focus on the CP-even two-bottom operators, including the dipole operators (assuming real Wilson coefficients) and those that modify the SM $Zb\bar{b}$ couplings. We consider the rate (the total cross section, or $R_{b}$) and forward-backward asymmetry measurements, using the projected measurement reaches in the recent Snowmass2021 studies CEPCPhysicsStudyGroup:2022uwl ; deBlas:2022ofj . Since the Z-pole measurements alone are not sufficient to simultaneously constrain all four parameters, a high energy run (e.g. at 240 GeV) becomes crucial for the global analysis. Furthermore, as we will show, there is non-trivial interplay between the linear contributions of dim-6 operators and the dipole-squared contributions. As a result, the fitting results with the Z-pole and 240 GeV measurements exhibit a nontrivial second minimum due to cancellations from linear and quadratic contributions. An additional run at a different energy (e.g. at 360 GeV) helps to resolve this second minimum.
@@
We also note that the bottom forward-backward asymmetry ($A^{0,b}_{\rm FB}$) measurement at LEP exhibits a $2.5\sigma$ deviation from the SM prediction ALEPH:2005ab . Many studies have been carried out on the possible new physics explanations and future experimental tests of them (see e.g. Refs. Choudhury:2001hs ; He:2003qv ; DaRold:2010as ; Gori:2015nqa ; Liu:2017xmc ; Yan:2021veo ; Yan:2021htf ; Breso-Pla:2021qoe ; Chai:2022ued ; Dong:2022ayy ; Yan:2023ccj ; Bishara:2023qhe ). We find that a non-zero bottom dipole interaction also provides a possible explanation of the $A^{0,b}_{\rm FB}$ discrepancy. In particular, the dipole-squared contribution tends to reduce the asymmetry which is in the direction that the LEP measurement prefers. Both the LEP $A^{0,b}_{\rm FB}$ discrepancy and the bottom dipole explanation can be thoroughly tested at a future lepton collider.
@@
The rest of this paper is organized as follows: In Section II, we introduce the SMEFT framework as well as the effective couplings for our analysis. The details of the analysis are described in Section III. Our results in terms of the global fit with four effective couplings are presented in Section IV, including the preferred regions in the parameter space and the numerical values of the one-sigma bounds and correlations. In Section V, we study the impacts of dipole operators on the LEP $A^{0,b}_{\rm FB}$ discrepancy. Finally, we summarize our work in Section VI. In Appendix A, we provide the results in terms of the dim-6 Wilson coefficients.
@@ II The SMEFT framework
where $q_{L}\equiv\begin{pmatrix}t_{L}\\ b_{L}\end{pmatrix}$ is the third generation quark doublet. [^2] In a more general framework, these operators can have more complex flavor structures. Here we consider only the third generation quarks. The dipole operators are not hermitian and can have complex coefficients in general. Here we focus on CP-even effects and assume the dipole coefficients are real. After electroweak symmetry breaking (EWSB), we have the following interactions between the bottom quark and the EW gauge bosons
@@
where the new physics contribution $\delta g_{Lb}$ and $\delta g_{Rb}$ are related to the coefficients of the first three operators in Eq. (missing) 2 (the convention of the Wilson coefficients follows the one in Eq. (missing) 1)
@@
The choice of EW input scheme have little impact on our study since we focus on the bottom measurements. On the other hand, the choice of the bottom mass has a significant impact since it determines the size of the interference term between SM and the dipole contribution. We use the value in the $\overline{\rm MS}$ scheme with the renormalization scale $\mu=m_{b}$, which gives $m_{b}\approx 4.18$ GeV ParticleDataGroup:2024cfk . The reason to choose $\mu=m_{b}$ instead of $\mu=m_{Z}$ is that here the bottom mass characterize the helicity flip of the bottom quark, so that a definition closer to the pole mass is more appropriate. This is different from the treatment of the bottom mass in the $h\to b\bar{b}$ process, which needs to be run to $\mu=m_{h}$ since it is proportional to the bottom Yukawa coupling.
@@ III Measurements and Analyses
Given that we focus on the bottom operators, we also consider only the observables that are directly related to these operators. At the Z-pole, the main observables are $R_{b}$, $A^{0,b}_{\rm FB}$ and $A_{b}$, which are summarized in Table 1. The current measured values (from LEP and SLD) and the SM predictions are from the PDG ParticleDataGroup:2024cfk (see also Ref. ALEPH:2005ab ), while the projected CEPC reach are from Ref. CEPCPhysicsStudyGroup:2022uwl . Note that, without beam polarization, CEPC could not directly measure $A_{b}$. However, the projection of $A_{b}$ is provided in Ref. CEPCPhysicsStudyGroup:2022uwl , which is derived from the measurements of $A^{0,b}_{\rm FB}$ and $A_{e}$, while the projection of $A^{0,b}_{\rm FB}$ is not provided. We therefore directly use the projection of $A_{b}$ reported in Ref. CEPCPhysicsStudyGroup:2022uwl . For the measurements at higher energies, there is no official CEPC projections. We use the projections in Ref. deBlas:2022ofj on the total cross section of $e^{+}e^{-}\to b\bar{b}$ and $A^{0,b}_{\rm FB}$ at 240240240\,240GeV and 360360360\,360GeV which are based on simple estimations with only statistical uncertainties. In particular, an overall universal efficiency of $15\%$ is applied on the signal events, while $\cos\theta$ (where $\theta$ is the production polar angle) is assumed to be within the range [0.9, 0.9]0.90.9[-0.9,\,0.9][ - 0.9 , 0.9 ]. The projections are summarized in Table 2. While the WW threshold run is potentially relevant for our study, there is no corresponding projections from Ref. deBlas:2022ofj , and we do not include it in the analysis. For all future measurements, we assume the central values are SM-like.
@@
We compute all observables at the tree level in terms of the four effective parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ described in the previous section. To be consistent with the SM predictions in Table 1 which include loop corrections, the SMEFT predictions are multiplied by the ratio between SM full prediction and SM tree-level prediction, when necessary. Note this procedure is most relevant for the analysis with current measurements in Section V. Two scenarios are considered: In the first one, the observables are expanded in terms of the effective parameters and only the linear contributions are kept. In the second one, the full expression is used without any expansion. [^3] This case does not necessarily corresponds to a quadratic relation since $R_{b}$, $A^{0,b}_{\rm FB}$ and $A_{b}$ are ratios in which the effective parameters contribute to both the numerator and the denominator. Nevertheless, the numerical difference is small beyond the square terms.

### v2

@@ Abstract
While often omitted in the SMEFT analyses of electroweak measurements, the electroweak dipole operators of the bottom quark have been found to be important in some cases and are also related to processes involving the top quark. In this paper, we further investigate their effects, focusing on the measurements of the $e^{+}e^{-}\to b\bar{b}$ process at a future lepton collider. Their linear contributions are suppressed by the bottom mass due to the helicity flip in the interference term with the SM amplitude, leading to a nontrivial interplay between the linear and the quadratic contributions. With two runs at the Z-pole and 240 GeV, the effects of CP-even dipole coefficients can be well separated from the modifications of the SM $Zb\bar{b}$ couplings, while an additional run (e.g. at 360 GeV) is useful for lifting a nontrivial second best-fit point due to the quadratic contributions.
@@ I Introduction
The Standard Model (SM) of particle physics has been tested by collider experiments for several decades. In the electroweak (EW) sector, the SM predictions are in good agreement with the precision measurements carried out at the Large Electron-Positron collider (LEP) and Stanford Linear Collider (SLC) experiments ALEPH:2005ab ; ALEPH:2013dgf , as well as those at the Large Hadron Collider (LHC). Higgs measurements have been performed in multiple channels at the LHC, and good precision has already been achieved in some channels with the accumulated data ATLAS:2022vkf ; CMS:2022dwd . Several future lepton collider projects have also been proposed, including CEPC CEPCStudyGroup:2018ghi ; CEPCPhysicsStudyGroup:2022uwl , FCC-ee FCC:2018evy ; Bernardi:2022hny , ILC ILCInternationalDevelopmentTeam:2022izu , C3 Bai:2021rdg , CLIC CLICdp:2018cto ; CLIC:2018fvx , as well as a possible high-energy muon collider Aime:2022flm ; Accettura:2023ked . These future projects hold the potential to significantly enhance the current measurement precision of Higgs and EW observables, providing an even better probe of the SM and the possible new physics beyond it.
Assuming the scale of the potential new physics is significantly larger than the EW scale, the Standard Model Effective Field Theory (SMEFT) provides a systematic and model-independent parameterization of the new physics effects, where deviations from the SM predictions (from new physics effects) are given by the Wilson coefficients of higher dimensional operators,
@@
where $d$ is the dimension of the operator and $\Lambda$ parameterizes the scale of the potential underlying new physics. The dimension-six operators are expected to constitute the leading effects to the experimental observables (besides the dimension-5 operator which generates Majorana neutrino masses) and have been thoroughly studied. Many global analyses have been performed, with both the current measurements and the projected future ones Falkowski:2015jaa ; Durieux:2017rsg ; Barklow:2017suo ; Durieux:2018tev ; Durieux:2018ggn ; Ellis:2018gqa ; DeBlas:2019qco ; Durieux:2019rbz ; Ellis:2020unq ; Bruggisser:2022rhb ; Liu:2022vgo ; deBlas:2022ofj ; Ethier:2021bye ; Brivio:2022hrb ; Bartocci:2023nvp ; Allwicher:2023shc ; Grunwald:2023nli ; Garosi:2023yxg ; Elmer:2023wtr ; Kassabov:2023hbm ; Bellafronte:2023amz ; Celada:2024mcf ; Asteriadis:2024qim . In many cases, good overall precision can be achieved despite the large number of free parameters — a clear demonstration of the usefulness of the global SMEFT framework.
@@
While the top dipole operators are considered in many SMEFT analyses at high-energy colliders Buckley:2015lku ; Schulze:2016qas ; CMS:2019too ; ATLAS:2019fwo ; CMS:2021klw , the dipole operators of other fermions are usually omitted, and for good reasons. [^1] On the other hand, see Refs. daSilvaAlmeida:2019cbr ; Wen:2024cfu on studies of light-quark dipoles at colliders and Ref. Wen:2023xxc on probing electron dipole operators with transverse spin asymmetry. Refs. Biswas:2021qaf ; Biswas:2022fsr also discussed the relevant dipole operators in the VBF Higgs and Higgs-photon production processes at the LHC. First, the SM amplitude and the amplitude with one dipole insertion have different helicity structures, and a mass insertion is needed to obtain an interference term. For light fermions, this means that the new physics effects are strongly suppressed by fermion mass at the $1/\Lambda^{2}$ level and mainly given by the dipole-squared terms, which are at the $1/\Lambda^{4}$ level. Second, low-energy electric or magnetic dipole measurements provide very stringent constraints on some of the dipole operators (especially the electron and muon ones Muong-2:2006rrc ; ACME:2018yjb ; Muong-2:2008ebm ; Muong-2:2021ojo ) which are far beyond the reaches of collider experiments. However, for the second-heaviest fermion, the bottom quark, the above two reasons are not fully applicable, and the effects of dipole operators are indeed found to be important in some cases. For instance, in Ref. Liu:2022vgo , it was found that the linear contributions of bottom dipole in EW observables, while suppressed by the bottom mass, tend to be more important than the loop effects of top operators, and the former should not be ignored if the latter is included in the analysis. On the other hand, Ref. Durieux:2019rbz studied the quadratic contributions of bottom dipole operators (while setting the bottom mass to zero) and found their effects to be important in a global framework. Together, these two studies suggest that both the linear and the quadratic contributions of the bottom dipole operators could be important, and an analysis including the full contributions of bottom dipole operators is needed to understand their impacts.
@@
In this paper, we conduct a further investigation into the effects of bottom dipole operators in the $e^{+}e^{-}\to b\bar{b}$ process, with their full contributions (both linear and quadratic) taken into account. We focus on the scenario of a future circular lepton collider, with a tera-Z program that significantly improves the EW measurements at LEP/SLD, as well as runs at both 240 GeV and around the $t\bar{t}$ threshold (around 360 GeV). As a first step towards a more comprehensive global analysis, in this study, we focus on the CP-even two-bottom operators, including the dipole operators (assuming real Wilson coefficients) and those that modify the SM $Zb\bar{b}$ couplings. We consider the rate (the total cross section, or $R_{b}$) and forward-backward asymmetry measurements, using the projected measurement reaches in recent Snowmass 2021 studies CEPCPhysicsStudyGroup:2022uwl ; deBlas:2022ofj . Since the Z-pole measurements alone are not sufficient to simultaneously constrain all four parameters, a high energy run (e.g. at 240 GeV) becomes crucial for the global analysis. Furthermore, as we will show, there is nontrivial interplay between the linear contributions of dim-6 operators and the dipole-squared contributions. As a result, the fitting results with the Z-pole and 240 GeV measurements exhibit a nontrivial second minimum due to cancellations from linear and quadratic contributions. An additional run at a different energy (e.g. at 360 GeV) helps to resolve this second minimum.
@@
We also note that the bottom forward-backward asymmetry ($A^{0,b}_{\rm FB}$) measurement at LEP exhibits a $2.5\sigma$ deviation from the SM prediction ALEPH:2005ab . Many studies have been carried out on the possible new physics explanations and future experimental tests of them (see e.g. Refs. Choudhury:2001hs ; He:2003qv ; DaRold:2010as ; Gori:2015nqa ; Liu:2017xmc ; Yan:2021veo ; Yan:2021htf ; Breso-Pla:2021qoe ; Chai:2022ued ; Dong:2022ayy ; Yan:2023ccj ; Bishara:2023qhe ). A non-zero bottom dipole interaction would also modify the SM prediction for $A^{0,b}_{\rm FB}$. In particular, the dipole-squared contribution tends to reduce the asymmetry which is in the direction that the LEP measurement prefers. However, we find that a large coefficient ($|\kappa_{bZ}|\sim 0.001$, defined in Eq. (missing) 3) is needed for a good fit with the LEP measurement. Assuming dipole interactions are generated at the one-loop level, this would require the new physics scale to be very low, likely out of the perturbative SMEFT regime.
@@
The rest of this paper is organized as follows. In Section II, we introduce the SMEFT framework as well as the effective couplings for our analysis. The details of the analysis are described in Section III. Our results in terms of the global fit with four effective couplings are presented in Section IV, including the preferred regions in the parameter space and the numerical values of the 1-sigma bounds and correlations. Finally, we summarize our work in Section V. In Appendix A, we provide the results in terms of the dim-6 Wilson coefficients. In Appendix B, we provide additional results which use a different bottom mass value (with the renormalization scale $\mu=m_{Z}$ instead of $\mu=m_{b}$).
@@ II The SMEFT framework
where $q_{L}\equiv\begin{pmatrix}t_{L}\\ b_{L}\end{pmatrix}$ is the third generation quark doublet. [^2] In a more general framework, these operators can have more complex flavor structures. Here we consider only the third generation quarks. The dipole operators are not hermitian and can have complex coefficients in general. Here we focus on CP-even effects and assume that the dipole coefficients are real. The imaginary parts of dipole operators, which give raises to CP-violation effects, are stringently constrained by the measurements of the electron electric dipole moment (EDM). For instance, Ref. Panico:2018hal considered the two-loop effects on the electron EDM in SMEFT and found that the size of the coefficients Im[$c_{bW}$] and Im[$c_{bB}$] are constrained to be less than one, assuming $\Lambda=10\,\text{TeV}$, based on the measurement by the ACME collaboration in 2018 ACME:2018yjb . These constraints are likely to be further improved by the next generation ACME experiment.
After electroweak symmetry breaking (EWSB), we have the following interactions between the bottom quark and the EW gauge bosons
@@
where the new physics contributions $\delta g_{Lb}$ and $\delta g_{Rb}$ are related to the coefficients of the first three operators in Eq. (missing) 2 (the convention of the Wilson coefficients follows the one in Eq. (missing) 1)
@@
The choice of EW input scheme has little impact on our study, since we focus on the bottom measurements. On the other hand, the choice of the bottom mass has a significant impact, since it determines the size of the interference term between the SM and the dipole contribution. We use the value in the $\overline{\rm MS}$ scheme with the renormalization scale $\mu=m_{b}$, which gives $m_{b}\approx 4.18$ GeV ParticleDataGroup:2024cfk . The reason to choose $\mu=m_{b}$ instead of $\mu=m_{Z}$ is that here the bottom mass characterizes the helicity flip of the bottom quark, so that a definition closer to the pole mass is more appropriate. This is different from the treatment of the bottom mass in the $h\to b\bar{b}$ process, which needs to be run to $\mu=m_{h}$.
For the latter, the running of the bottom mass (or more precisely, the bottom Yukawa coupling) effectively resums the leading logs of the higher-order QCD corrections, which is important for obserables such as $\Gamma(h\to b\bar{b})$. On the other hand, the observables $R_{b}$ and $A^{0,b}_{\rm FB}$ are ratios of hadronic observables where one might expect large cancellations of QCD corrections. A proper treatment of the higher-order QCD corrections is also beyond the scope of our study. For comparison, we also show in Appendix B the fit results that use $\mu=m_{Z}$ instead of $\mu=m_{b}$. We note that the choice of bottom mass has also been discussed in previous studies (see e.g. Refs. Catani:1999nf ; Bernreuther:2016ccf ; deBlas:2024bmz ) but in the context of SM predictions for the Z-pole observables (such as $A^{0,b}_{\rm FB}$), for which the impacts are relatively small.
@@ III Measurements and Analyses
Given that we focus on the bottom operators, we also consider only the observables that are directly related to these operators. At the Z-pole, the main observables are $R_{b}$, $A^{0,b}_{\rm FB}$ and $A_{b}$, which are summarized in Table 1.333While other observables, such as the total hadronic cross section at Z-pole, $\sigma_{had}^{0}$, also receive contributions from the bottom operators, they are also sensitive to many other operators, and are thus not included in our analysis. The current measured values (from LEP and SLD) and the SM predictions are from the PDG ParticleDataGroup:2024cfk (see also Ref. ALEPH:2005ab ), while the projected CEPC reach are from Ref. CEPCPhysicsStudyGroup:2022uwl . Note that, without beam polarization, CEPC could not directly measure $A_{b}$. However, the projection of $A_{b}$ is provided in Ref. CEPCPhysicsStudyGroup:2022uwl , which is derived from the measurements of $A^{0,b}_{\rm FB}$ and $A_{e}$, while the projection of $A^{0,b}_{\rm FB}$ is not provided. We therefore directly use the projection of $A_{b}$ reported in Ref. CEPCPhysicsStudyGroup:2022uwl . For measurements at higher energies, there are no official CEPC projections. We use the projections in Ref. deBlas:2022ofj on the total cross section of $e^{+}e^{-}\to b\bar{b}$ and $A^{0,b}_{\rm FB}$ at 240 GeV and 360 GeV which are based on simple estimations with only statistical uncertainties. In particular, an overall universal efficiency of $15\%$ is applied to signal events, while $\cos\theta$ (where $\theta$ is the production polar angle) is assumed to be within the range [0.9, 0.9]0.90.9[-0.9,\,0.9][ - 0.9 , 0.9 ]. The projections are summarized in Table 2. While the WW threshold run is potentially relevant for our study, there is no corresponding projections from Ref. deBlas:2022ofj , and we do not include it in the analysis. For all future measurements, we assume that the central values are SM-like.
@@
We compute all observables at the tree level in terms of the four effective parameters $\left\{\delta g_{Lb},\,\delta g_{Rb},\,\kappa_{bA},\,\kappa_{bZ}\right\}$ described in the previous section. To be consistent with the SM predictions in Table 1 which include loop corrections, the SMEFT predictions are multiplied by the ratio between SM full prediction and SM tree-level prediction, when necessary. Two scenarios are considered: In the first one, the observables are expanded in terms of the effective parameters and only the linear contributions are kept. In the second one, the full expression is used without any expansion. [^4] This case does not necessarily corresponds to a quadratic relation since $R_{b}$, $A^{0,b}_{\rm FB}$ and $A_{b}$ are ratios in which the effective parameters contribute to both the numerator and the denominator. Nevertheless, the numerical difference is small beyond the square terms.